TheJag18:

12th grade math (Mathematics for Data and Financial Literacy) A medical school student opens a credit card to pay for living expenses during their first year of medical school. The student expects to pay $12,500 in living expenses. The credit card company charges an annual interest rate of 15.99% compounded continuously. If no payments are made for three years, what will be the balance on the card rounded to the nearest penny? $20,194.87 $20,129.60 $18,496.25 $14,667.42

Phantomdex:

Give me a moment I'm not the best at math but let me see if I can solve this.

TheJag18:

its 12th grade

Phantomdex:

Hm, based on my math and a calculator I use to help me, I believe the answer should be A. $20,194.87. This should be because if we take the amount of money we start with and compound it continuously by the percentage. First, convert R as a percent to r as a decimal r = R/100 r = 15.99/100 r = 0.1599 rate per year, Then solve the equation for A A = Pert A = 12,500.00(2.71828)(0.1599)(3) A = $20,194.87 Summary: The total amount accrued, principal plus interest, with compound interest on a principal of $12,500.00 at a rate of 15.99% per year compounded continuously over 3 years is $20,194.87.

jayfafr:

20,186.76.

hannahj:

$18,496.25 :)

Phantomdex:

Yeah I'm not the best at math :']

TheJag18:

its a

hannahj:

sorry I did that wrong it is A. mb :/

TheJag18:

i already said its a

Phantomdex:

:0 So I was right!

hannahj:

@thejag18 wrote:
i already said its a
ok chill jeez

TheJag18:

yeah you got it right and hannah, i did nothing to warrent you saying chill, i was just saying i said its a.

TheJag18:

An arithmetic sequence begins with 25, 31, 37, 43, 49 … Which option below represents the formula for the sequence? f(n) = 25 + 6(n) f(n) = 25 + 6(n + 1) f(n) = 25 + 6(n − 1) f(n) = 19 + 6(n + 1)

Phantomdex:

@thejag18 wrote:
An arithmetic sequence begins with 25, 31, 37, 43, 49 … Which option below represents the formula for the sequence? f(n) = 25 + 6(n) f(n) = 25 + 6(n + 1) f(n) = 25 + 6(n − 1) f(n) = 19 + 6(n + 1)
I believe that the answer should be f(n) = 25 + 6(n-1).

Phantomdex:

Since it seems to match the sequence the most.

TheJag18:

it is

TheJag18:

An investor deposited $7,590 in a savings account 10 years ago that earns simple annual interest. Today, the account balance is $11,916.30. What is the simple interest rate? 0.63% 6.3% 0.57% 5.7%

Phantomdex:

Sorry this one is taking me a moment, but please post new questions whenever you have one!

TheJag18:

oh ok, ill post on another one after this one

Phantomdex:

Based on the amount you've gotten I believe the answer would be D. 5.7%.